OpenStudy (anonymous):

sqrt of 1/3 plus sqrt of 48 plus fourth root of 144 minus 5 over sqrt of 3 ?

OpenStudy (anonymous):

\[\sqrt{1/3} +\sqrt{48}+\sqrt[4]{144}-5/\sqrt{3}=\]